Two packages are dropped from an airplane. A parachute can increase the cross-sectional area of each packages by a factor of 36. The parachute on package 1 fails to open, and the terminal speed of package 1 is 23.1 m/s. The parachute on package 2 opens. What is the terminal speed of package 2 in m/s? Enter a number with one digit after the decimal point.
Answer #1

Terminal velocity v = sqrt(2mg/C rho A)

v2/v1 = sqrt(A1/A2)

v2 / 23.1 = sqrt(1/36)

v2 = 23.1/6 = 3.85 m/s answer